# Exploring Dark Knowledge under Various Teacher Capacities and Addressing Capacity Mismatch

### v1

@@ IV-C The Relative Class Affinities
The second observation in Sect. IV-A implies that the relative probability values given by different teachers seem to be consistent. Hence, we explore this observation further through several quantitative metrics. The intuitive idea is to explore whether different teachers give the same top-$K$ predictions or not. Specifically, we train a smaller teacher $t_{\text{small}}$ and a larger teacher $t_{\text{large}}$ on the same dataset. Then, we calculate the predicted top-$K$ classes for the $i$-th training sample, and denote the predicted class sets as $\mathcal{C}_{i,K}^{t_{\text{large}}}$ and $\mathcal{C}_{i,K}^{t_{\text{small}}}$, respectively. We denote $|\mathcal{C}_{i,K}^{t_{\text{large}}}\cap\mathcal{C}_{i,K}^{t_{\text{small}}}|$ as the number of overlapped classes. Then we show the distribution of the number of overlapped classes when considering different $K$. Fig. 5 displays the results on CUB and TinyImageNet. The percentages in the figure show how many data samples have the number of overlapping classes corresponding to the values of the y-axis. Considering the top-$3$ predicted classes of ResNeXt50-32-4d and ResNeXt101-32-8d on CUB, there are $21.6\%$ training samples that have 3 common predicted classes, and $56.4\%$ training samples that have 2 common predicted classes. If we consider top-$8$ predicted classes, there are nearly $69.4\%$ training samples that have at least 5 common predicted classes. This implies that different teachers indeed have a great deal of consistency in top-$K$ class recognition.
@@
Figure 5: The predicted top-$K$ class set overlap of a smaller and a larger teacher network on two datasets. The x-axis shows $K$ and the y-axis shows the number of overlapped classes.
@@
Figure 6: Several magnitude-agnostic metrics that measure the consistency of relative class affinities between teacher networks with various capacities. The used dataset is CIFAR-10.
@@
where $j_{1}$ and $j_{2}$ show the index of classes. $C$ is the number of classes, and $\text{sg}(\cdot)$ returns 1 or 0 depending on whether the input is positive or not. These three metrics only depend on classes’ relative magnitudes and are irrelevant to the absolute values. The metrics are calculated on CIFAR-10 under a series of ResNeXt networks. We set $K=5$ for the rank set overlap metric. The above equations only show the metrics on a single training sample, and we report the average result on 50K training samples. The results are in Fig. 6. Excitingly, these metrics among teachers with different capacities do not vary a lot, and the Spearman correlations are almost all larger than $0.85$. According to the interpretation of Kendall’s $\tau$ [68, 69], if the smaller teacher predicts that class $j_{1}$ is more related to the target class than that of class $j_{2}$, then the larger teacher has a probability of $0.875$ to give the same relative affinity. These metrics demonstrate that teachers know approximately the same about relative class affinities. Certainly, the output logits of different networks are not completely identical in order, even for the same networks under different training randomness [70, 71].
IV-D Verification by Group Classification
@@
The probability vector $\mathbf{p}^{t}$ inherently represents the posterior probability distribution predicted by a teacher network. During the training process, the cross-entropy loss minimizes the divergence between the posterior label distribution and the empirical “one-hot” label distribution. The “one-hot” labels cannot introduce class relations into the learning process. Hence, the teacher’s label could be viewed as an alternate that could help the student grasp the affinities between classes. However, we do not know the real posterior distribution of each training sample and cannot judge which teacher’s label is better.
@@
Figure 7: The group classification on CIFAR-100. Each row shows the averaged posterior distribution given by a specific network. Circle size and color map determine the value. The three subfigures show results of fine-grained class “orchids”, “hamster”, and “maple”, respectively, which belong to the superclass “flowers”, “small mammals”, and “trees”. The x-axis shows the $3\times 5=15$ fine-grained classes in the three superclasses.
@@
Although it is hard to assign the real posterior distribution for each training sample, the group classification scenes could provide some prior information between categories. CIFAR-100 contains 100 categories, which contain 20 superclasses and 5 classes for each superclass  [^1] https://www.cs.toronto.edu/~kriz/cifar.html. For example, the superclass “flowers” contains five fine-grained classes including “orchids”, “poppies”, “roses”, “sunflowers” and “tulips”. One hypothesis is that categories in the same superclass are more similar to each other. We train different teachers on CIFAR-100 and find that the hypothesis holds true for most superclasses. Fig. 7 shows the averaged posterior distribution on three superclasses, i.e., “flowers”, “small mammals”, and “trees”. For each superclass, we select its first fine-grained class as the target label, i.e., “orchids”, “hamster”, and “maple”. For each target label, we obtain its corresponding training samples, i.e., 500 samples for each. We calculate the averaged probabilities via $\mathbb{E}_{\mathbf{x},y=c}[\mathbf{p}^{t}(\mathbf{x};\tau)]$ for $c$ being one of the three target classes. We set $\tau$ as $4.0$. The teacher network ranges from ResNet14 to ResNeXt29-64-4d. The circle size and color map represent the probability value. Obviously, for a specific fine-grained class, the predicted probability values for the classes in the same group are larger than other classes. Additionally, the relative probability values for the classes in the same group are basically consistent among the investigated networks. For example, different teacher networks agree that fine-grained classes “oak” and “willow” are more relevant to the target category “maple”, while “palm” is less relevant. Although teachers are consistent in the relative class affinities, the variance of the probability values becomes smaller as the teacher capacity increases. For example, the difference in circle sizes in the last row is much smaller than that in the first row, which behaves similarly in each subfigure.
@@
Then we provide some rules that define the similarity between classes. For example, the rule “$c_{1}$: $c_{2}$, $c_{3}$, $c_{4}$, $c_{5}$” represents that a sample in class $c_{1}$ is predicted with the following rank: $\mathbf{p}_{c_{1}}>\mathbf{p}_{c_{2}}>\mathbf{p}_{c_{3}}>\mathbf{p}_{c_{4}}>\mathbf{p}_{c_{5}}$. If a network gives the rank as “$c_{1}$, $c_{3}$, $c_{2}$, $c_{4}$, $c_{5}$”, then the Kendall’s $\tau$ between the prediction and the rule is 0.8. To measure the absolute difference in the predictions of non-ground-truth classes, we also calculate the standard deviation of the probabilities, i.e., $\sigma\left(\left[\mathbf{p}_{c_{2}},\mathbf{p}_{c_{3}},\mathbf{p}_{c_{4}},\mathbf{p}_{c_{5}}\right]\right)$. The following lists the rules defined in CIFAR-100:
@@
The above rules are concluded from the averaged posterior probability distributions shown in Fig. 7. Notably, we only consider the class relationship in the same superclass for simplification. Additionally, we also construct a group classification task that utilizes five digits classification datasets, i.e., MNIST, MNISTM, USPS, SVHN, SYN. We denote this as Digits-Five, which is often used in domain adaptation tasks [38] or task relationship estimation [72]. The construction process is simple where we only centralize the data together and obtain a task with $5\times 10=50$ classes. According to the transfer performances and the task relation coefficients presented in the previous works [38, 72], we abstract several rules as follows:
@@
Although it is difficult to claim that the above rules apply to all samples of the corresponding class, the rules may provide a potentially approximate estimation for the real posterior of a specific sample. Tab. I lists the results of Kendall’s $\tau$ and the standard deviation of non-ground-truth classes’ probabilities. The consistency degrees of teachers’ outputs with the defined rules do not vary obviously. The predicted ranks are almost completely consistent in the simpler classification task, i.e., Kendall’s $\tau$ is almost near 1.0 in Digits Five. However, the standard deviations of larger teachers are much smaller than those of smaller teachers, implying that larger teachers’ outputs are less varied between non-ground-truth classes. These experimental studies on group classification verify the initial observations in Sect. IV-A again.
TABLE I: The consistency between the ranks predicted by different teachers with the defined rules and the standard deviation of the non-ground-truth class probability values.
@@
V Addressing Capacity Mismatch
As introduced in Sect. III, the “capacity mismatch” refers to the phenomenon that larger networks may not teach students as well as smaller teachers [33, 32, 35, 36, 31]. The above observations show that the relative class affinity between the larger teacher and smaller teacher’s outputs are basically consistent, while the absolute probabilities between non-ground-truth classes are not so discriminative for larger teachers. Hence, our proposed methods aim to enhance the distinctness between non-ground-truth class probabilities from different perspectives.
V-A Fusing Global Class Relations (FGCR)
. For each class $c$, we calculate the averaged probabilities via $\mathbb{E}_{\mathbf{x},y=c}[\mathbf{p}(\mathbf{x};\tau_{0})]$, and advocate that this captures the global relationships among classes. According to Fig. 7, the global relationships given by different teachers are similar in relative values, which may be useful for stable distillation. Hence, we fuse this to each training sample of the $c$-th class as follows:
@@ V-A Fusing Global Class Relations (FGCR)
where $\alpha$ is the hyper-parameter. We empirically found that setting $\tau_{0}<\tau$ for calculating the averaged probabilities could enhance the variance of non-ground-truth class probabilities. As shown in Fig. 7, the averaged probability distribution provided by larger networks still shows less varied values among non-ground-truth classes, which may limit the effectiveness of this method. Hence, we only provide this method as a trial, and it is not the focus of our paper. In fact, the performances of FGCR could surpass the distillation performance of a smaller teacher, while it is still worse than other advanced techniques as shown in Tab. II.
V-B Regularizing Teachers (RegT)
. During the phase of training teachers, we add the following regularization to enhance the variance of non-ground-truth class probabilities:
@@ V-B Regularizing Teachers (RegT)
where $\mathbf{q}(\mathbf{x};1.0)=[\mathbf{p}_{c}(\mathbf{x};1.0)]_{c\neq y}$ and $\beta$ is the coefficient of the regularization term. The loss function explicitly encourages the teacher network to decrease the influence of the ground-truth class and enhance the distinctness of non-ground-truth classes. Different from other post-processing methods, this way could enhance the variance of non-ground-truth class probabilities in the stage of training teachers.
V-C Instance-Specific Asymmetric Temperature Scaling (ISATS)
. The previous work [31] proposes the approach named asymmetric temperature scaling (ATS) to tackle the capacity mismatch in KD. ATS applies different temperatures to the logits of ground-truth and non-ground-truth classes, i.e.,
@@ V-C Instance-Specific Asymmetric Temperature Scaling (ISATS)
TABLE II: Performance comparisons of addressing capacity mismatch in KD. The last three rows are our methods.
@@
where $\mathbf{q}(\mathbf{x};\tau)$ denotes the probability vector of non-ground-truth classes, i.e., $\mathbf{q}(\mathbf{x};\tau)=[\mathbf{p}_{c}(\mathbf{x};\tau)]_{c\neq y}$ and $\mathbf{p}(\mathbf{x};\tau)=\text{SF}(\mathbf{f}(\mathbf{x});\tau)$. $v(\cdot)$ calculates the variance of the elements in a vector. That is, the optimal temperature for each instance is searched to make the probabilities of non-ground-truth classes more distinct. Then, we set $\tau_{1}(\mathbf{x})=\tau^{\star}+1$ and $\tau_{2}(\mathbf{x})=\tau^{\star}$ as recommended by ATS [31]. Compared with ATS, the proposed ISATS has two advantages. First, ISATS saves the effort of finding the proper value of $\tau_{2}$, and thus the setting of $\tau_{1}$ could also be quickly determined. Second, ISATS considers a more fine-grained way to release the discriminative information contained in non-ground-truth classes for each training sample. Therefore, ISATS could commonly obtain better performances than ATS.
VI Performance Comparisons
This section provides performance comparisons of addressing capacity mismatch in KD. The experimental studies follow the settings in [31] and utilize the publicly available code  [^2] https://github.com/lxcnju/ATS-LargeKD. We directly cite some experimental results for comparisons. Specifically, the utilized datasets are CIFAR-10/CIFAR-100 [37], TinyImageNet [59], CUB [57], and Stanford Dogs [56]. Teacher networks include versions of ResNet [54], WideResNet [58], and ResNeXt [55]. Student networks are VGG [61], ShuffleNetV1/V2 [4, 5], AlexNet [73], and MobileNetV2 [3]. We train networks on corresponding datasets for 240 epochs. SGD optimizer with a momentum value of 0.9 is used. The learning rate is 0.05 by default, and the batch size is 128. For our proposed FGCR, we set $\tau_{0}=\tau-1$ and $\alpha\in\{0.1,0.5\}$, and the best results are reported. For RegT, $\beta$ is searched in $\{0.01,0.001\}$. For ISATS, we search for the best temperature for each instance in the scope of $\{1.0,2.0,3.0,4.0,5.0,6.0,8.0\}$. The distribution of the optimal instance-specific temperatures on CIFAR-100 is as follows: 4.0 accounts for about $31\%$, 5.0 accounts for about $27\%$, 3.0 accounts for about $22\%$, and others account for $20\%$.
TABLE III: Ensemble performance comparisons of addressing capacity mismatch in KD. The last three rows are our methods.

### v2

@@ 4.3 The Relative Class Affinities
The second observation in Sect. 4.1 implies that the relative probability values given by different teachers seem to be consistent. Hence, we explore this observation further through several quantitative metrics. The intuitive idea is to explore whether different teachers give the same top-$K$ predictions or not. Specifically, we train a smaller teacher $t_{\text{small}}$ and a larger teacher $t_{\text{large}}$ on the same dataset. Then, we calculate the predicted top-$K$ classes for the $i$-th training sample, and denote the predicted class sets as $\mathcal{C}_{i,K}^{t_{\text{large}}}$ and $\mathcal{C}_{i,K}^{t_{\text{small}}}$, respectively. We denote $|\mathcal{C}_{i,K}^{t_{\text{large}}}\cap\mathcal{C}_{i,K}^{t_{\text{small}}}|$ as the number of overlapped classes. Then we show the distribution of the number of overlapped classes when considering different $K$. Fig. 5 displays the results on CUB and TinyImageNet. The percentages in the figure show how many data samples have the number of overlapping classes corresponding to the values of the y-axis. Considering the top-$3$ predicted classes of ResNeXt50-32-4d and ResNeXt101-32-8d on CUB, there are $21.6\%$ training samples that have 3 common predicted classes, and $56.4\%$ training samples that have 2 common predicted classes. If we consider top-$8$ predicted classes, there are nearly $69.4\%$ training samples that have at least 5 common predicted classes. This implies that different teachers indeed have a great deal of consistency in top-$K$ class recognition.
@@
Figure 5: The predicted top-$K$ class set overlap of a smaller and a larger teacher network on two datasets. (Left) CUB dataset; (Right) TinyImageNet dataset
@@
Figure 6: Several magnitude-agnostic metrics that measure the consistency of relative class affinities between teacher networks with various capacities. (Left) Rank Set Overlap Ratio; (Middle) Spearman Correlation; (Right) Kendall’s Tau
@@
where $j_{1}$ and $j_{2}$ show the index of classes. $C$ is the number of classes, and $\text{sg}(\cdot)$ returns 1 or -1 depending on whether the input is positive or not. These three metrics only depend on classes’ relative magnitudes and are irrelevant to the absolute values. The metrics are calculated on CIFAR-10 under a series of ResNeXt networks. We set $K=5$ for the rank set overlap metric. The above equations only show the metrics on a single training sample, and we report the average result on 50K training samples. The results are in Fig. 6. Excitingly, these metrics among teachers with different capacities do not vary a lot, and the Spearman correlations are almost all larger than $0.85$. According to the interpretation of Kendall’s $\tau$ [71, 72], if the smaller teacher predicts that class $j_{1}$ is more related to the target class than that of class $j_{2}$, then the larger teacher has a probability of $0.875$ to give the same relative affinity. Certainly, the output logits of different networks are not completely identical in order, even for the same networks under different training randomness [73, 74]. These metrics demonstrate that teachers know approximately the same about relative class affinities.
4.4 Verification by Group Classification
@@
During training, the cross-entropy loss minimizes the divergence between the predicted posterior distribution and the empirical one-hot label distribution. Although one-hot labels do not capture class relationships, the posterior distribution does. The teacher network’s probability vector, $\mathbf{p}^{t}$, inherently represents this distribution, serving as an alternative that may help the student learn class affinities. However, since the true posterior distribution for each sample is unknown, it is difficult to ascertain which teacher’s label is superior.
@@
Figure 7: The group classification on CIFAR-100. Each row shows the averaged posterior distribution given by a specific network. Circle size and color map determine the value. The three subfigures show results of fine-grained class “orchids", “hamster", and “maple", respectively, which belong to the superclass “flowers", “small mammals", and “trees". The x-axis shows the $3\times 5=15$ fine-grained classes in the three superclasses.
@@
While precise posterior distribution assignment per training sample remains challenging, group classification scenarios offer valuable inter-category priors. CIFAR-100 contains 100 categories, which contain 20 superclasses and 5 classes for each superclass  [^1] https://www.cs.toronto.edu/~kriz/cifar.html. For example, the superclass “flowers" contains five fine-grained classes including “orchids", “poppies", “roses", “sunflowers" and “tulips". One hypothesis is that categories in the same superclass are more similar to each other. We train different teachers (ranges from ResNet14 to ResNeXt29-64-4d) on CIFAR-100 and find that the hypothesis holds true for most superclasses. Fig. 7 presents averaged posterior distribution on three superclasses: “flowers", “small mammals", and “trees". For each superclass, we select its first fine-grained class as the target label, i.e., “orchids", “hamster", and “maple". For each target label, we obtain its corresponding training samples, i.e., 500 samples for each. We calculate the averaged probabilities via $\mathbb{E}_{\mathbf{x},y=c}[\mathbf{p}^{t}(\mathbf{x};\tau)]$ ($\tau$=4.0) for $c$ being one of the three target classes. The circle size and color map represent the probability value. Key observations: (1) Target classes exhibit stronger affinity to intra-superclass members across all networks; (2) Relative intra-group probabilities maintain cross-architectural consistency. For example, different teacher networks agree that fine-grained classes “oak" and “willow" are more relevant to the target category “maple", while “palm" is less relevant. (3) Probability variance decreases with increasing model capacity, evident from diminishing circle size differences in deeper networks. For example, the difference in circle sizes in the last row is much smaller than that in the first row, which behaves similarly in each subfigure.
@@
Then we provide some rules that define the similarity between classes. For example, the rule “$c_{1}$: $c_{2}$, $c_{3}$, $c_{4}$, $c_{5}$" represents that a sample in class $c_{1}$ is predicted with the following rank: $\mathbf{p}_{c_{1}}>\mathbf{p}_{c_{2}}>\mathbf{p}_{c_{3}}>\mathbf{p}_{c_{4}}>\mathbf{p}_{c_{5}}$. If a network gives the rank as “$c_{1}$, $c_{3}$, $c_{2}$, $c_{4}$, $c_{5}$", then the Kendall’s $\tau$ between the prediction and the rule is 0.8. To measure the absolute difference in the predictions of non-ground-truth classes, we also calculate the standard deviation of the probabilities, i.e., $\sigma\left(\left[\mathbf{p}_{c_{2}},\mathbf{p}_{c_{3}},\mathbf{p}_{c_{4}},\mathbf{p}_{c_{5}}\right]\right)$. The following lists the rules defined in CIFAR-100:
@@
These rules are derived from the averaged posterior probability distributions in Fig. 7. For simplicity, we only consider intra-superclass relationships. We further construct a group classification task using five digit recognition datasets: MNIST, MNIST-M, USPS, SVHN, and SYN. We denote this as Digits-Five, which is often used in domain adaptation tasks [38] or task relationship estimation [75]. The construction process is simple where we only centralize the data together and obtain a task with $5\times 10=50$ classes. Based on transfer performance analyses and task relation coefficients from existing research [38, 75], we abstract several rules as follows:
@@
While these rules may not apply to every sample, they may provide an approximate estimation of a sample’s true posterior distribution. Table 1 presents the results for Kendall’s $\tau$ and the standard deviation of probabilities for non-ground-truth classes. The consistency of teacher outputs with the defined rules shows little variation; in the simpler Digits Five task, the predicted ranks are nearly identical (with Kendall’s $\tau$ approaching 1.0). However, larger teachers yield much smaller standard deviations than smaller ones, indicating less variability among non-ground-truth classes. These experiments on group classification further validate the initial observations in Sect. 4.1.
Table 1: The consistency between the ranks predicted by different teachers with the defined rules and the standard deviation of the non-ground-truth class probability values.
@@
5 Addressing Capacity Mismatch
As introduced in Sect. 3, the “capacity mismatch” refers to the phenomenon that larger networks may not teach students as well as smaller teachers [33, 32, 35, 36, 31]. The above observations show that the relative class affinity between the larger teacher and smaller teacher’s outputs are basically consistent, while the absolute probabilities between non-ground-truth classes are not so discriminative for larger teachers. Hence, our proposed methods aim to enhance the distinctness between non-ground-truth class probabilities from different perspectives. The relative class affinity we observed ensures the correctness and effectiveness of our methods.
5.1 Fusing Global Class Relations (FGCR)
For each class $c$, we calculate the averaged probabilities via $\mathbb{E}_{\mathbf{x},y=c}[\mathbf{p}(\mathbf{x};\tau_{0})]$, and advocate that this captures the global relationships among classes. According to Fig. 7, the global relationships given by different teachers are similar in relative values, which may be useful for stable distillation. Hence, we fuse this to each training sample of the $c$-th class as follows:
@@ 5.1 Fusing Global Class Relations (FGCR)
where $\alpha$ is the hyper-parameter. The relative class affinity suggests that small teacher can be regarded as a form of label smoothing (LS) for the large teacher, as both ensure consistency in the rank order of classes’ probabilities. Currently, the large teacher exhibits an overly one-hot output, necessitating LS to increase the variance of non-ground-truth class probabilities. Although LS can hurt knowledge KD [46],  [76] demonstrates that label smoothing is compatible with KD at lower temperatures. Therefore, we set $\tau_{0}<\tau$. As shown in Fig. 7, the averaged probability distribution provided by larger networks still shows less varied values among non-ground-truth classes, which may limit the effectiveness of this method. Hence, we only provide this method as a trial, and it is not the focus of our paper. In fact, the performances of FGCR could surpass the distillation performance of a smaller teacher, while it is still worse than other advanced techniques as shown in Tab. 2.
5.2 Regularizing Teachers (RegT)
During the phase of training teachers, we add the following regularization to enhance the variance of non-ground-truth class probabilities:
@@ 5.2 Regularizing Teachers (RegT)
where $\mathbf{q}(\mathbf{x};1.0)=[\mathbf{p}_{c}(\mathbf{x};1.0)]_{c\neq y}$ and $\beta$ is the coefficient of the regularization term. Equation (8) explicitly encourages teacher network to decrease the influence of the ground-truth class and enhance the distinctness of non-ground-truth classes. Different from other post-processing methods, this way could enhance the variance of non-ground-truth class probabilities in the stage of training teachers, which could provide an excellent initialization for distillation.
5.3 Instance-Specific Asymmetric Temperature Scaling (ISATS)
The previous work [31] proposes the approach named asymmetric temperature scaling (ATS) to tackle the capacity mismatch in KD. ATS applies different temperatures to the logits of ground-truth and non-ground-truth classes, i.e.,
@@ 5.3 Instance-Specific Asymmetric Temperature Scaling (ISATS)
Table 2: Performance comparisons of addressing capacity mismatch in KD. The last three rows are our methods.
@@
where $\mathbf{q}(\mathbf{x};\tau)$ denotes the probability vector of non-ground-truth classes in teacher network, i.e., $\mathbf{q}(\mathbf{x};\tau)=[\mathbf{p}_{c}(\mathbf{x};\tau)]_{c\neq y}$ and $\mathbf{p}(\mathbf{x};\tau)=\text{SF}(\mathbf{f}(\mathbf{x});\tau)$. It is noted that, according to relative class affinity, enlarging the teacher model does not alter the rank order among non-ground-truth classes. Consequently, all non-target classes share the same $\tau$, eliminating the need to apply different $\tau$ values across classes. $v(\cdot)$ calculates the variance of the elements in a vector. That is, the optimal temperature for each instance is searched to make the probabilities of non-ground-truth classes more distinct. Then, we set $\tau_{1}(\mathbf{x})=\tau^{\star}+1$ and $\tau_{2}(\mathbf{x})=\tau^{\star}$ as recommended by ATS [31]. Equation (10) has no closed-form solution. So like ATS, ISATS still requires manual search to ensure an appropriate temperature. Nevertheless, ISATS has two advantages over ATS. First, ISATS saves the effort of finding the proper value of $\tau_{2}$ and $\tau_{1}$. ATS requires performing multiple distillation processes and comparing the experimental results at different temperatures to determine proper $\tau$. In contrast, ISATS can directly compute $\tau^{\star}$ based on (10) in advance, eliminating the need for comparative distillation experiments. Second, ISATS considers a more fine-grained way to release the discriminative information contained in wrong classes for each training sample. In summary, ISATS is an all-around improvement over ATS.
6 Performance Comparisons
This section provides performance comparisons of addressing capacity mismatch in KD. The experimental studies follow the settings in [31] and utilize the publicly available code  [^2] https://github.com/lxcnju/ATS-LargeKD. We directly cite some experimental results for comparisons. Specifically, the utilized datasets are CIFAR-10/CIFAR-100 [37], TinyImageNet [62], CUB [60], and Stanford Dogs [59]. We choose teacher models significantly larger than their corresponding student models as we do in Section 4. Teacher networks include versions of ResNet [57], WideResNet [61], and ResNeXt [58]. Student networks are VGG [64], ShuffleNetV1/V2 [4, 5], AlexNet [77], and MobileNetV2 [3]. We train networks on corresponding datasets for 240 epochs. SGD optimizer with a momentum value of 0.9 is used. The learning rate is 0.05 by default, and the batch size is 128. For our proposed FGCR, we set $\tau_{0}=\tau-1$ and $\alpha\in\{0.1,0.5\}$, and the best results are reported. For RegT, $\beta$ is searched in $\{0.01,0.001\}$. For ISATS, we search for the best temperature for each instance in the scope of $\{1.0,2.0,3.0,4.0,5.0,6.0,8.0\}$. The distribution of the optimal instance-specific temperatures on CIFAR-100 is as follows: 4.0 accounts for about $31\%$, 5.0 accounts for about $27\%$, 3.0 accounts for about $22\%$, and others account for $20\%$.
Table 3: Ensemble performance comparisons of addressing capacity mismatch in KD. The last three rows are our methods.